OpenStudy (anonymous):

find all angles theta between 0 and 180 satisfying the equation cos theta = 1/9

OpenStudy (anonymous):

try \(\cos^{-1}(\frac{1}{9})\) with a calculator

OpenStudy (anonymous):

\[\cos \theta=\frac{ 1 }{ 9 },\theta =\cos^{-1} \frac{ 1 }{9 }=83.62 degree\]